OpenStudy (anonymous):

I will give you a MEDAL if you get this right! A pharmacist is to prepare 15 milliliters of special eye drops for a glaucoma patient. The eye-drop solution must have a 4% active ingredient, but the pharmacist only has 10% solution and 1% solution in stock. How much of each type of solution should be used to fill the prescription?

jimthompson5910 (jim_thompson5910):

x+y = 15 y = 15 - x ------------------ 0.10x + 0.01y = 0.04*15 0.10x + 0.01y = 0.6 0.10x + 0.01(15 - x) = 0.6 0.10x + 0.01(15) - 0.01(x) = 0.6 0.10x + 0.15 - 0.01x = 0.6 0.09x + 0.15 = 0.6 0.09x = 0.6 - 0.15 0.09x = 0.45 x = 0.45/0.09 x = 5

jimthompson5910 (jim_thompson5910):

now use this to find y y = 15 - x y = 15 - 5 y = 10

jimthompson5910 (jim_thompson5910):

So the pharmacist needs 5 milliliters of the 10% solution and 10 milliliters of the 1% solution.

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

np